OpenStudy (anonymous):

The following table shows the values of f(x) and g(x) for different values of x: x 1 2 3 4 5 f(x) 5 25 125 625 3.125 g(x) -3 -3 -3 -3 -3 Which of the following best describes the function f(x) + g(x)? 5x – 3x –3(5x) 5x – 3 5x – 3

OpenStudy (inkyvoyd):

dat pic, going to get you in trouble with preetha >.>

OpenStudy (anonymous):

@superdavesuper

OpenStudy (anonymous):

g(x) is always -3 f(x) should be 5^x

OpenStudy (anonymous):

so A ?

OpenStudy (anonymous):

a is 5^x-3x

OpenStudy (anonymous):

i would go with \(5^x-3\) as well

OpenStudy (anonymous):

or c? 5^x-3

OpenStudy (anonymous):

if that is an option

OpenStudy (anonymous):

yeah, C it is always C

OpenStudy (anonymous):

yea it is, its C, ok that makes more sense

OpenStudy (anonymous):

yeah seems you misstyped the answers in your question

OpenStudy (the_fizicx99):

Well, f(x) is being squared, 5^2 = 25 5^3 = 125 So, 5^x - 3

OpenStudy (anonymous):

ok how about this question ? The table below shows the average temperature, in degrees Celsius, in Jacob's city over a period of five months: Month 1 2 3 4 5 Temperature 5 7.2 9.4 11.6 13.8 Did the temperature in Jacob's city increase linearly or exponentially? Linearly, because the table shows a constant percentage increase in temperature each month. Exponentially, because the table shows a constant percentage increase in temperature each month. Linearly, because the table shows that temperature increased by the same amount each month. Exponentially, because the table shows that temperature increased by the same amount each month.

OpenStudy (anonymous):

i think its D

OpenStudy (the_fizicx99):

Well, the average rate of change is 2.2, 5 + 2.2 = 7.2 7.2 + 2.2 = 9.4 9.4 + 2.2 = 11.6 So clearly its linear

OpenStudy (anonymous):

it increases by 2.2 every time so c ?

OpenStudy (the_fizicx99):

Ya

OpenStudy (anonymous):

ok that was easier than it seemed